OpenStudy (anonymous):

Can someone explain to me in steps how to do, Solve for x: (x+4)/3 = 2?

OpenStudy (lalaly):

\[\cancel{3} \times \frac{ x+4}{\cancel{3}}= 2\times 3\]\[x+4=6\]\[x=6-4=2\]

OpenStudy (anonymous):

Lol, I wanted to beat you to it. But, I didn't! :D

OpenStudy (lalaly):

lol:D

OpenStudy (anonymous):

nice use of \cancel

OpenStudy (lalaly):

i learn from the best :D